# Shallow IQP circuits and graph generation

### v1

@@ II Graphs and Random Graph Distributions
A graph distribution defines a probability distribution over the space of graphs satisfying a given criterion. In this work, we restrict attention to undirected, unweighted graphs to isolate structural dependencies while maintaining a compact binary representation. Such graphs connect nodes by undirected edges: if node $i$ is linked to node $j$, then node $j$ is equally linked to node ii. Consequently, the adjacency matrix $\mathbf{A}$ is symmetric, $A_{ij}=A_{ji}$, and self-loops are excluded, i.e. $A_{ii}=0\ \forall\ i$.
@@ III IQP circuit born machines
Our objective is to train shallow IQP models to generate graphs according to a ground truth distribution pp. Specifically, we aim to learn a parameterized distribution $q_{\bm{\theta}}$ such that, for any feature of interest $f(\mathbf{z})$,
@@ III.1 Shallow IQP circuits
A shallow IQP circuit on NN qubits is defined by the following sequential operations,
@@ III.2 Optimization via the maximum mean discrepancy
for vectors $\bm{x},\bm{y}\in\{0,1\}^{n}$. Since each bit contributes either 0 or 11 to the sum, the squared Euclidean distance $\|\bm{x}-\bm{y}\|^{2}$ coincides with the Hamming distance, i.e., the number of positions at which the entries of the vectors differ. This equivalence allows the MMD to be expressed in terms of Pauli-Z observables [rudolphTrainabilityBarriersOpportunities2024],
@@
Table 1: Summary of the datasets. $\bar{\rho}$ denotes the average density of the samples, BP (%) the fraction of bipartite graphs, $\bar{\beta}$ the average spectral bipartivity, and NN the total number of samples.
@@ IV.4 Sampling
After training, we generate samples by preparing the NN-qubit state with the optimized parameters on the IBM Aachen quantum computer and measuring in the computational basis. Each measurement produces an NN-bit string $\mathbf{z}$, which we map to a graph following the procedure in Sec. II.1.

### v2

@@ II Graphs and Random Graph Distributions
A graph distribution defines a probability distribution over the space of graphs satisfying a given criterion. In this work, we restrict attention to undirected, unweighted graphs to isolate structural dependencies while maintaining a compact binary representation. Such graphs connect nodes by undirected edges: if node $i$ is linked to node $j$, then node $j$ is equally linked to node $i$. Consequently, the adjacency matrix $\mathbf{A}$ is symmetric, $A_{ij}=A_{ji}$, and self-loops are excluded, i.e. $A_{ii}=0\ \forall\ i$.
@@ III IQP circuit born machines
Our objective is to train shallow IQP models to generate graphs according to a ground truth distribution $p$. Specifically, we aim to learn a parameterized distribution $q_{\bm{\theta}}$ such that, for any feature of interest $f(\mathbf{z})$,
@@ III.1 Shallow IQP circuits
A shallow IQP circuit on $N$ qubits is defined by the following sequential operations,
@@ III.2 Optimization via the maximum mean discrepancy
for vectors $\bm{x},\bm{y}\in\{0,1\}^{n}$. Since each bit contributes either 0 or $1$ to the sum, the squared Euclidean distance $\|\bm{x}-\bm{y}\|^{2}$ coincides with the Hamming distance, i.e., the number of positions at which the entries of the vectors differ. This equivalence allows the MMD to be expressed in terms of Pauli-Z observables [rudolphTrainabilityBarriersOpportunities2024],
@@
Table 1: Summary of the datasets. $\bar{\rho}$ denotes the average density of the samples, BP (%) the fraction of bipartite graphs, $\bar{\beta}$ the average spectral bipartivity, and $N$ the total number of samples.
@@ IV.4 Sampling
After training, we generate samples by preparing the $N$-qubit state with the optimized parameters on the IBM Aachen quantum computer and measuring in the computational basis. Each measurement produces an $N$-bit string $\mathbf{z}$, which we map to a graph following the procedure in Sec. II.1.